OpenStudy (anonymous):

How do you find the "greatest common factor" between the numbers 21, 30, 44?

OpenStudy (saifoo.khan):

1

OpenStudy (anonymous):

Easiest is to write up the prime factors: 21=3*7 30=2*3*5 44=2*2*11 Now there is nothing that appears in all so gcd(21,30,44)=1